OpenStudy (ineedhelponhw):

If an orbiting satellite has a diameter of 60 cm what is the total surface area of the satellite? surface area = 4pier^2

JoelTheBoss (joel_the_boss):

What do know about this question? Is there anything in general you need help on?

OpenStudy (ineedhelponhw):

I forgot how to do it. Its been years

OpenStudy (ineedhelponhw):

And its the only question I have so I can get my other homework done.

JoelTheBoss (joel_the_boss):

\[4*\pi(30*10^{-2})^2\] This is the equation correct?

OpenStudy (ineedhelponhw):

Ye it is I did what i could remeber

JoelTheBoss (joel_the_boss):

ok would you like me to solve it for you? Step by step?

OpenStudy (ineedhelponhw):

Yes please so I can get other questions like that.

OpenStudy (ineedhelponhw):

I can solve bymyself

OpenStudy (ineedhelponhw):

laster on

JoelTheBoss (joel_the_boss):

What are asking for exactly? Another equation similar to this one?

OpenStudy (ineedhelponhw):

They want the total surface area of the satellite which you know has a diameter of 60cm

OpenStudy (ineedhelponhw):

And they gave us the hint of the surface area 4pir^2

JoelTheBoss (joel_the_boss):

ok then you solve the equation I gave you to get the answer. :)

OpenStudy (ineedhelponhw):

O

JoelTheBoss (joel_the_boss):

Would you like me to do it? Or are you ok..

OpenStudy (ineedhelponhw):

May you please.

OpenStudy (ineedhelponhw):

I want to know step by step so I can later on.

JoelTheBoss (joel_the_boss):

typo

OpenStudy (ineedhelponhw):

o ok

JoelTheBoss (joel_the_boss):

First we can simplify the equation to look like \[4*(\pi(30*10^{-2}))\]

OpenStudy (ineedhelponhw):

Okay.

OpenStudy (ineedhelponhw):

We have to simplify diameter right? so that's why it's 30

JoelTheBoss (joel_the_boss):

The next step would be to write \[30*10^{-2}\]In Proper Scientific Notation. \[4*(\pi(3*10^{-1}))\]

OpenStudy (ineedhelponhw):

Okay.

JoelTheBoss (joel_the_boss):

Next we convert \( 30*10^{-1}\) from scientific notation \(4*(pi*0.3)\)

JoelTheBoss (joel_the_boss):

Then we Multiply \(pi\) by 0.03 to get \[0.9424778\]Next just plug the 4 back in, \(4*0.9424778\)

JoelTheBoss (joel_the_boss):

Can you do the rest or do you want me to do it?

OpenStudy (ineedhelponhw):

You do it please

JoelTheBoss (joel_the_boss):

ok :)

JoelTheBoss (joel_the_boss):

The Last step is just to times \(4*0.9424778\) The final answer would be, \(3.76991118\) Hope I helped. :)

OpenStudy (ineedhelponhw):

Thank you so much!

JoelTheBoss (joel_the_boss):

Any time :)